(Everything in series)


What frequency should be applied to an RLC circuit with R=30 ?, C=1.5 pF, and L=0.01 ?H in order to achieve the largest maximum current?
Answer #1

largest max current will occur at resonant frequency

omega=1/sqrt(L*C)=8.2*10^7

Frequency=omega/(2*pi)=1.3*10^7
Answer #2

Resonant frequency occurs at    \(\omega_0=1/\sqrt{LC}\)   

and

    \(\omega =2\pi f\)   

combining gives:

1/((2pi)sqrt(LC)) = 1.3x10^7 Hz